# Crossed products by compact group actions with the weak tracial Rokhlin property

### v2

@@ 5. Permanence properties
Let $F\subseteq A\otimes_{min}B$ and $S\subseteq C(G)$ be finite sets, let $\varepsilon>0$, $x\in(A\otimes_{min}B)_{+}$ with $\|x\|=1$, and $y\in(A\otimes_{min}B)^{\alpha\otimes\beta}_{+}\setminus\{0\}$. We need to find a contractive positive element $d\in(A\otimes_{min}B)^{\alpha\otimes\beta}$ and a completely positive contractive map $\psi\colon C(G)\to\overline{d(A\otimes_{min}B)d}$ such that
@@
We may assume there exist $c_{1},c_{2},\dots,c_{m}\in A$ and $e_{1},e_{2},\dots,e_{m}\in B$ such that $F=\{c_{i}\otimes e_{i}\colon i=1,2,\dots,m\}$ and $\|c_{i}\|\leq 1,\|e_{i}\|\leq 1$ for $i=1,2,\dots,m$. Choose $\delta$ such that $0<\delta<\frac{\varepsilon}{3}$ and $(1-5\delta-\delta^{2})>1-\varepsilon$. Also, choose $\delta_{1}$ such that $\frac{1}{2}<\delta_{1}<1$ and $(1-5\delta-\delta^{2})\delta_{1}^{2}>1-\varepsilon$. Set $z=(x-\delta_{1})_{+}$. Since $\beta$ is trivial, $(A\otimes_{min}B)^{\alpha\otimes\beta}\cong A^{\alpha}\otimes_{min}B^{\beta}$. By Kirchberg’s Slice Lemma (see Lemma 2.8), there exist $x_{1}\in A_{+}$ and $x_{2}\in B_{+}$ such that $x_{1}\otimes x_{2}\precsim_{A\otimes_{min}B}z$ with $\|x_{1}\|=\|x_{2}\|=1$, and $y_{1}\in(A^{\alpha})_{+}$ and $y_{2}\in(B^{\beta})_{+}$ such that $y_{1}\otimes y_{2}\precsim_{(A\otimes_{min}B)^{\alpha\otimes\beta}}y$ with $\|y_{1}\|=\|y_{2}\|=1$. By [32, Lemma 2.2, Lemma 2.4(i)], there exists $w\in A\otimes_{min}B$ such that $\|wxw^{*}-x_{1}\otimes x_{2}\|<\delta^{2}$ and $\|w\|\leq\delta_{1}^{-\frac{1}{2}}<\delta_{1}^{-1}$. Thus there exists $v\in A\otimes_{min}B$, where $v=\sum_{i=1}^{k}v_{i}\otimes w_{i}$ for some $v_{i}\in A$ and $w_{i}\in B$, $i=1,2,\dots,k$ such that
@@
Set $E=\{c_{i}\colon i=1,2,\dots,m\}\cup\{v_{i}\colon i=1,2,\dots,k\}$. Since $B$ and $B^{\beta}$ are simple and unital, by [31, Proposition 2.7(v)], there exist $n,m\in\mathbb{N}$ such that
@@
Set $M=1+\sum_{i=1}^{k}\|v_{i}\|+\sum_{i=1}^{k}\|w_{i}\|$. Choose $\eta>0$ such that $\eta<\frac{\delta}{Mk}$. Applying Lemma 3.7 to $\alpha$ for $\eta$, $E$, $S$, $x_{0}$, $x_{1}$, and $y_{1}$, we obtain a contractive positive element $r\in A^{\alpha}$ and a completely positive contractive map $\varphi\colon C(G)\to\overline{rAr}$ such that
@@
Assuming (1), by [37, Lemma 4.12], we have $b\in A$ such that $1=b^{*}ab\precsim_{A}a$.
For the converse, assuming (2), by parts (1) and (8) of Lemma 2.5, we have $b\in A$ such that $1=b^{*}ab$. By [37, Lemma 4.12], AA is purely infinite. ∎
@@
By Lemma 5.3, we need to show that for any $a\in(A^{\alpha})_{+}\setminus\{0\}$, we have $1\precsim_{A^{\alpha}}a$. We may assume $\|a\|=1$. Since $A^{\alpha}$ is simple and non-type I, by Lemma 2.6, there exist two nonzero mutually orthogonal Cuntz equivalent contractive positive elements $a_{0},a_{1}$ in $A^{\alpha}$ with $\|a_{0}\|=1$ and $\|a_{1}\|=1$ such that $a_{0}+a_{1}\precsim_{A^{\alpha}}a$.
@@
Set $F=\{a_{1}\}$. For any $n\in\mathbb{N}$, suppose $\{S_{n}\}$ is a sequence of finite sets in $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $S_{n}$, $F_{2}$, $x$, and $y$, there exist a contractive positive element $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Define a completely positive contractive order zero map $\varphi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\varphi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. Denote $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. We have
@@
Since $A$ is simple and purely infinite, by [37, Lemma 4.12], there exists $x\in A$ such that $x^{*}a_{1}x=1$. Therefore, by Theorem 2.16 and Corollary 2.18, we have
@@
By Theorem 5.4, the algebra $A^{\alpha}$ is purely infinite. Since $A^{\alpha}$ and $A\rtimes_{\alpha}G$ are stably isomorphic (see Corollary 4.9), by [31, Theorem 4.23], $A\rtimes_{\alpha}G$ is also purely infinite. ∎
@@
[28, Definition 2.1] Let $A$ be a unital C*-algebra with $A\ncong\mathbb{C}$. Then $A$ is tracially $\mathcal{Z}$-absorbing (or tracially $\mathcal{Z}$-stable) if and only if for any finite set $F\subseteq A$, $\varepsilon>0$, $x\in A_{+}\setminus\{0\}$, and $n\in\mathbb{N}$, there exists a completely positive contractive order zero map $\psi\colon M_{n}\to A$ such that
@@
[37, Proposition 4.16] Let $n\in\mathbb{N}$ and let $\varepsilon>0$. Then there exists $\delta>0$ satisfying the following property. Let $D$ be a C*-algebra and for $j,k=1,2,\dots,n$ let $y_{j,k}\in D$, satisfying (in the unitization of $D$)
@@
We need to show that for any $\varepsilon>0$, any finite set $F\subseteq A^{\alpha}$, any $x\in(A^{\alpha})_{+}\setminus\{0\}$, and any $n\in\mathbb{N}$, there exists a completely positive contractive order zero map $\psi\colon M_{n}\to A^{\alpha}$ such that
@@
Since $A$ is tracially $\mathcal{Z}$-stable, for given $n$ and $F$, for $\varepsilon_{0}$ and for $x_{1}$, we have a completely positive contractive order zero map $\psi_{0}\colon M_{n}\rightarrow A$ such that
@@
Suppose $\{S_{n}\}$ is a sequence of finite sets in $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F_{0}\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $S_{n}$, $F\cup\{x_{1}\}$, $x_{2}$, there exists a contractive positive element $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Define a completely positive contractive order zero map $\varphi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\varphi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. Denote $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. We have
@@
By Corollary 2.21, we have $r-\varphi\circ\psi_{0}(1)=\varphi(1-\psi_{0}(1))\precsim_{(A^{\alpha})_{\infty}}\varphi(x_{1})=r^{\frac{1}{2}}x_{1}r^{\frac{1}{2}}\precsim_{(A^{\alpha})_{\infty}}x_{1}$. By part (1) of Lemma 2.5, for $\frac{\varepsilon}{64}$ there exists $\delta_{1}>0$ such that $(\varphi(1-\psi_{0}(1))-\frac{\varepsilon}{64})_{+}\precsim_{(A^{\alpha})_{\infty}}(x_{1}-\delta_{1})_{+}$. Therefore, by part (8) of Lemma 2.5, there exists $v\in(A^{\alpha})_{\infty}$ such that
@@
Thus we can choose $u=(u_{n})_{n\in\mathbb{N}}\in l^{\infty}(\mathbb{N},A^{\alpha})$ such that $\pi_{A^{\alpha}}(u)=v$ and a completely positive contractive map $\varphi_{n_{0}}\colon A\to\overline{d_{n_{0}}A^{\alpha}d_{n_{0}}}$ such that
@@
By [37, Lemma 4.17], we obtain a completely positive contractive order zero map $\psi\colon M_{n}\to\overline{d_{n_{0}}A^{\alpha}d_{n_{0}}}$ such that
@@
The algebra $A$ is $\mathcal{Z}$-stable, hence by [28, Proposition 2.2] it is tracially $\mathcal{Z}$-stable. Therefore, by Theorem 5.9, $A^{\alpha}$ is tracially $\mathcal{Z}$-stable. $A$ is amenable, so $A^{\alpha}$ is also amenable; thus by [28, Theorem 4.1], $A^{\alpha}$ is $\mathcal{Z}$-stable. By [51, Corollary 3.2], $\mathcal{Z}$-stability is preserved under stable isomorphism, and $A^{\alpha}$ and $A\rtimes_{\alpha}G$ are stably isomorphic (see Corollary 4.9), so the crossed product $A\rtimes_{\alpha}G$ is also $\mathcal{Z}$-stable. ∎
@@
We first prove that if $B$ is a unital C*-algebra, $a,g\in B$ satisfy $0\leq a,g\leq 1$, and let $\varepsilon_{1},\varepsilon_{2}>0$. Then
@@
Indeed, using Lemma 2.7 in the first step, part (10) of Lemma 2.5 in the second step, and part (2) of 2.5 in the third step, we get
@@
Set $F=\{a,b,a^{\prime},b^{\prime}\}$. Suppose $\{S_{n}\}$ is a sequence of finite sets in $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $S_{n}$, $F$, $f(b)$, there exists a contractive positive element $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Define a completely positive contractive order zero map $\varphi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\varphi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. Denote $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. We have
@@
So we can choose $u=(u_{n})_{n\in\mathbb{N}}\in l^{\infty}(\mathbb{N},A^{\alpha})$ such that $\pi_{A^{\alpha}}(u)=v$ and a completely positive contractive map $\varphi_{n_{0}}\colon A\to\overline{d_{n_{0}}A^{\alpha}d_{n_{0}}}$ such that
@@
The following proposition is a generalization of [37, Proposition 4.22]. This result is weaker than its counterpart for compact group actions with the Rokhlin property; namely, we cannot obtain an embedding of the Cuntz semigroup of the fixed point algebra into the Cuntz semigroup of the original algebra, but only an embedding between the subsemigroups consisting of purely positive elements (i.e., elements not Cuntz equivalent to a projection).
@@
The proof is essentially the same as that of [37, Proposition 4.22] when we use Proposition 5.12 in place of [37, Proposition 4.21], so we omit it. ∎
@@
Let $\tau^{\prime}\in\mathrm{QT}(A)$. We will show that the map $\mathrm{QT}(\iota)\colon\mathrm{QT}(A)\to\mathrm{QT}(A^{\alpha})$, sending $\tau^{\prime}\mapsto\tau^{\prime}\circ\iota$, is surjective. Let $z\in(A^{\alpha})_{+}\setminus\{0\}$. Suppose $\{S_{n}\}$ is a sequence of finite sets in $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $S_{n}$, $F$, $z$, there exists a contractive positive element $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Define a completely positive contractive order zero map $\varphi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\varphi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. Denote $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. We have
@@
For any $n\in\mathbb{N}$, there exist a contractive positive element $r_{n}\in(A^{\alpha})_{\infty}\cap A^{\prime}$ and a completely positive contractive order zero map $\psi_{n}\colon A\to\overline{r_{n}(A^{\alpha})_{\infty}r_{n}}$, such that for each $v\in\mathrm{QT}((A^{\alpha})_{\infty})$, $d_{v}(1-r_{n})<\frac{1}{n}$. Now, let $\tau\in\mathrm{QT}(A^{\alpha})$, which naturally induces a quasitrace on $(A^{\alpha})_{\infty}$, still denoted by $\tau$. We consider the weak-* limit of the quasitraces $\tau_{\infty}\circ\psi_{n}$ in $\mathrm{QT}(A)$, denoted $w^{*}-\lim(\tau_{\infty}\circ\psi_{n})$. Note that for $a\in A_{+}$ with $\|a\|\leq 1$ and $v\in\mathrm{QT}((A^{\alpha})_{\infty})$, $v(a)\leq d_{v}(a)$. Then for each $a\in A^{\alpha}$
@@
The proof is essentially the same as that of [7, Theorem 4.1] when we use Proposition 5.12 in place of [7, Lemma 3.7], so we omit it. ∎
@@
The proof is essentially the same as that of [48, Theorem 4.4], so we omit it. ∎
@@
The proof is essentially the same as that of [48, Theorem 4.6], so we omit it. ∎

### v3

@@ 5. Permanence properties
Let $F\subseteq A\otimes_{min}B$ and $S\subseteq C(G)$ be finite sets, let $\varepsilon>0$, $x\in(A\otimes_{min}B)_{+}$ with $\|x\|=1$, and $y\in(A\otimes_{min}B)^{\alpha\otimes\beta}_{+}\setminus\{0\}$. We need to find a positive contraction $d\in(A\otimes_{min}B)^{\alpha\otimes\beta}$ and a completely positive contractive map $\psi\colon C(G)\to\overline{d(A\otimes_{min}B)d}$ such that the following hold:
@@
We may assume there exist $c_{1},c_{2},\dots,c_{m}\in A$ and $e_{1},e_{2},\dots,e_{m}\in B$ such that $F=\{c_{i}\otimes e_{i}\colon i=1,2,\dots,m\}$ and $\|c_{i}\|\leq 1,\|e_{i}\|\leq 1$ for $i=1,2,\dots,m$. Choose $\delta$ such that $0<\delta<\frac{\varepsilon}{3}$ and $(1-5\delta-\delta^{2})>1-\varepsilon$. Also, choose $\delta_{1}$ such that $\frac{1}{2}<\delta_{1}<1$ and $(1-5\delta-\delta^{2})\delta_{1}^{2}>1-\varepsilon$. Set $z=(x-\delta_{1})_{+}$. Since $\beta$ is trivial, $(A\otimes_{min}B)^{\alpha\otimes\beta}\cong A^{\alpha}\otimes_{min}B^{\beta}$. By Kirchberg’s Slice Lemma (see Lemma 2.8), there exist $x_{1}\in A_{+}$ and $x_{2}\in B_{+}$ such that $x_{1}\otimes x_{2}\precsim_{A\otimes_{min}B}z$ with $\|x_{1}\|=\|x_{2}\|=1$, and $y_{1}\in(A^{\alpha})_{+}$ and $y_{2}\in(B^{\beta})_{+}$ such that $y_{1}\otimes y_{2}\precsim_{(A\otimes_{min}B)^{\alpha\otimes\beta}}y$ with $\|y_{1}\|=\|y_{2}\|=1$. By [31, Lemma 2.2, Lemma 2.4(i)], there exists $w\in A\otimes_{min}B$ such that $\|wxw^{*}-x_{1}\otimes x_{2}\|<\delta^{2}$ and $\|w\|\leq\delta_{1}^{-\frac{1}{2}}<\delta_{1}^{-1}$. Thus there exists $v\in A\otimes_{min}B$, where $v=\sum_{i=1}^{k}v_{i}\otimes w_{i}$ for some $v_{i}\in A$ and $w_{i}\in B$, $i=1,2,\dots,k$ such that
@@
Set $E=\{c_{i}\colon i=1,2,\dots,m\}\cup\{v_{i}\colon i=1,2,\dots,k\}$. Since $B$ and $B^{\beta}$ are simple and unital, by [30, Proposition 2.7(v)], there exist $n,m\in\mathbb{N}$ such that
@@
Set $M=1+\sum_{i=1}^{k}\|v_{i}\|+\sum_{i=1}^{k}\|w_{i}\|$. Choose $\eta>0$ such that $\eta<\frac{\delta}{Mk}$. Applying Lemma 3.7 to $\alpha$ for $\eta$, $E$, $S$, $x_{0}$, $x_{1}$, and $y_{1}$, we obtain a positive contraction $r\in A^{\alpha}$ and a completely positive contractive map $\varphi\colon C(G)\to\overline{rAr}$ such that the following hold:
@@
Assuming (1), by [36, Lemma 4.12], there exists $b\in A$ such that $1=b^{*}ab\precsim_{A}a$.
For the converse, assuming (2), by parts (1) and (8) of Lemma 2.5, we have $b\in A$ such that $1=b^{*}ab$. By [36, Lemma 4.12], AA is purely infinite. ∎
@@
By Lemma 5.3, we need to show that for any $a\in(A^{\alpha})_{+}\setminus\{0\}$, we have $1\precsim_{A^{\alpha}}a$. We may assume $\|a\|=1$. Since $A^{\alpha}$ is simple and non-type I, by Lemma 2.6, there exist two nonzero mutually orthogonal Cuntz equivalent positive contractions $a_{0},a_{1}$ in $A^{\alpha}$ with $\|a_{0}\|=1$ and $\|a_{1}\|=1$ such that $a_{0}+a_{1}\precsim_{A^{\alpha}}a$.
@@
Set $F=\{a_{1}\}$. For any $n\in\mathbb{N}$, suppose $\{S_{n}\}$ is a sequence of finite sets in $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $S_{n}$, $F_{2}$, $x$, and $y$, there exist a positive contraction $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that the following hold:
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Denote $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. Define a completely positive contractive order zero map $\varphi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\varphi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. We have the following hold:
@@
Since $A$ is simple and purely infinite, by [36, Lemma 4.12], there exists $x\in A$ such that $x^{*}a_{1}x=1$. Therefore, by Theorem 2.17 and Corollary 2.19, we have
@@
By Theorem 5.4, the algebra $A^{\alpha}$ is purely infinite. Since $A^{\alpha}$ and $A\rtimes_{\alpha}G$ are stably isomorphic (see Corollary 4.9), by [30, Theorem 4.23], $A\rtimes_{\alpha}G$ is also purely infinite. ∎
@@
[27, Definition 2.1] Let $A$ be a unital C*-algebra with $A\ncong\mathbb{C}$. Then $A$ is tracially $\mathcal{Z}$-absorbing (or tracially $\mathcal{Z}$-stable) if and only if for any finite set $F\subseteq A$, $\varepsilon>0$, $x\in A_{+}\setminus\{0\}$, and $n\in\mathbb{N}$, there exists a completely positive contractive order zero map $\psi\colon M_{n}\to A$ such that
@@
[36, Proposition 4.16] Let $n\in\mathbb{N}$ and let $\varepsilon>0$. Then there exists $\delta>0$ satisfying the following property. Let $D$ be a C*-algebra and for $j,k=1,2,\dots,n$, let $y_{j,k}\in D$, satisfying (in the unitization of $D$)
@@
We need to show that for any $\varepsilon>0$, any finite set $F\subseteq A^{\alpha}$, any $x\in(A^{\alpha})_{+}\setminus\{0\}$, and any $n\in\mathbb{N}$, there exists a completely positive contractive order zero map $\psi\colon M_{n}\to A^{\alpha}$ such that the following hold:
@@
Since $A$ is tracially $\mathcal{Z}$-stable, for given $n$ and $F$, for $\varepsilon_{0}$ and for $x_{1}$, we have a completely positive contractive order zero map $\psi_{0}\colon M_{n}\rightarrow A$ such that the following hold:
@@
Suppose $\{S_{n}\}$ is a sequence of finite sets in $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F_{0}\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $S_{n}$, $F\cup\{x_{1}\}$, $x_{2}$, there exists a positive contraction $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that the following hold:
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Denote $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. Define a completely positive contractive order zero map $\varphi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\varphi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. We have the following hold:
@@
By Corollary 2.22, we have $r-\varphi\circ\psi_{0}(1)=\varphi(1-\psi_{0}(1))\precsim_{(A^{\alpha})_{\infty}}\varphi(x_{1})=r^{\frac{1}{2}}x_{1}r^{\frac{1}{2}}\precsim_{(A^{\alpha})_{\infty}}x_{1}$. By part (1) of Lemma 2.5, for $\frac{\varepsilon}{64}$, there exists $\delta_{1}>0$ such that $(\varphi(1-\psi_{0}(1))-\frac{\varepsilon}{64})_{+}\precsim_{(A^{\alpha})_{\infty}}(x_{1}-\delta_{1})_{+}$. Therefore, by part (8) of Lemma 2.5, there exists $v\in(A^{\alpha})_{\infty}$ such that
@@
Thus we can choose $u=(u_{n})_{n\in\mathbb{N}}\in l^{\infty}(\mathbb{N},A^{\alpha})$ such that $\pi_{A^{\alpha}}(u)=v$ and a completely positive contractive map $\varphi_{n_{0}}\colon A\to\overline{d_{n_{0}}A^{\alpha}d_{n_{0}}}$ such that the following hold:
@@
By [36, Lemma 4.17], we obtain a completely positive contractive order zero map $\psi\colon M_{n}\to\overline{d_{n_{0}}A^{\alpha}d_{n_{0}}}$ such that
@@
The algebra $A$ is $\mathcal{Z}$-stable, hence by [27, Proposition 2.2] it is tracially $\mathcal{Z}$-stable. Therefore, by Theorem 5.9, $A^{\alpha}$ is tracially $\mathcal{Z}$-stable. $A$ is amenable, so $A^{\alpha}$ is also amenable; thus by [27, Theorem 4.1], $A^{\alpha}$ is $\mathcal{Z}$-stable. By [49, Corollary 3.2], $\mathcal{Z}$-stability is preserved under stable isomorphism, and $A^{\alpha}$ and $A\rtimes_{\alpha}G$ are stably isomorphic (see Corollary 4.9), so the crossed product $A\rtimes_{\alpha}G$ is also $\mathcal{Z}$-stable. ∎
@@
We first prove that if $B$ is a unital C*-algebra, $a,g\in B$ satisfy $0\leq a,g\leq 1$, and let $\varepsilon_{1},\varepsilon_{2}>0$, then
@@
Indeed, using Lemma 2.7 in the first step, part (10) of Lemma 2.5 in the second step, and part (2) of Lemma 2.5 in the third step, we get
@@
Set $F=\{a,b,a^{\prime},b^{\prime}\}$. Suppose $\{S_{n}\}$ is a sequence of finite sets in $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $S_{n}$, $F$, $f(b)$, there exist a positive contraction $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that the following hold:
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Denote $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. Define a completely positive contractive order zero map $\varphi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\varphi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. We have the following hold:
@@
So we can choose $u=(u_{n})_{n\in\mathbb{N}}\in l^{\infty}(\mathbb{N},A^{\alpha})$ such that $\pi_{A^{\alpha}}(u)=v$ and a completely positive contractive map $\varphi_{n_{0}}\colon A\to\overline{d_{n_{0}}A^{\alpha}d_{n_{0}}}$ such that the following hold:
@@
The following proposition is a generalization of [36, Proposition 4.22]. This result is weaker than its counterpart for compact group actions with the Rokhlin property; namely, we cannot obtain an embedding of the Cuntz semigroup of the fixed point algebra into the Cuntz semigroup of the original algebra, but only an embedding between the subsemigroups consisting of purely positive elements (i.e., elements not Cuntz equivalent to a projection).
@@
The proof is essentially the same as that of [36, Proposition 4.22] when we use Proposition 5.12 in place of [36, Proposition 4.21], so we omit it. ∎
@@
Let $\tau^{\prime}\in\mathrm{QT}(A)$. We will show that the map $\mathrm{QT}(\iota)\colon\mathrm{QT}(A)\to\mathrm{QT}(A^{\alpha})$, sending $\tau^{\prime}\mapsto\tau^{\prime}\circ\iota$, is surjective. Let $z\in(A^{\alpha})_{+}\setminus\{0\}$. Suppose $\{S_{n}\}$ is a sequence of finite sets in $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $S_{n}$, $F$, $z$, there exist a positive contraction $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that the following hold:
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Denote $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. Define a completely positive contractive order zero map $\varphi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\varphi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. We have the following hold:
@@
For any $n\in\mathbb{N}$, there exist a positive contraction $r_{n}\in(A^{\alpha})_{\infty}\cap A^{\prime}$ and a completely positive contractive order zero map $\psi_{n}\colon A\to\overline{r_{n}(A^{\alpha})_{\infty}r_{n}}$, such that for each $v\in\mathrm{QT}((A^{\alpha})_{\infty})$, $d_{v}(1-r_{n})<\frac{1}{n}$. Now, let $\tau\in\mathrm{QT}(A^{\alpha})$, which naturally induces a quasitrace on $(A^{\alpha})_{\infty}$, still denoted by $\tau$. We consider the weak-* limit of the quasitraces $\tau_{\infty}\circ\psi_{n}$ in $\mathrm{QT}(A)$, denoted $w^{*}-\lim(\tau_{\infty}\circ\psi_{n})$. Note that for $a\in A_{+}$ with $\|a\|\leq 1$ and $v\in\mathrm{QT}((A^{\alpha})_{\infty})$, $v(a)\leq d_{v}(a)$. Then for each $a\in A^{\alpha}$
@@
The proof is essentially the same as that of [6, Theorem 4.1] when we use Proposition 5.12 in place of [6, Lemma 3.7], so we omit it. ∎
@@
The proof is essentially the same as that of [47, Theorem 4.4], so we omit it. ∎
@@
The proof is essentially the same as that of [47, Theorem 4.6], so we omit it. ∎